At a work place, an employee is exposed to three chemicals during an eight hour shift as follows:

Chemical Actual Exposure (ppm)    PEL (ppm)
A 50 250
B 30 150
C 200     800    

What is the additive effects of multiple chemicals exposed during 8-hour period? _________

Is the worker required to wear a PPE?  ______
Answer #1

Additive effects of multiple chemicals = (C1÷L1+C2÷L2)+. . .(Cn÷Ln) =(50/250)+(30/150)+(200/800) =0.65

As additive effects of multiple chemicals is less than 1, so it is within the safe range and thus no need to wear a PPE.

KINDLY RATE THE ANSWER AS THUMBS UP. THANKS, A LOT